if C-C bond energy is 348 KJ/mol, C-H bond energy is 412 KJ/mol and H-H bond enery is 436 KJ/mol, the Pauling electronegative of C is about (Given electronegativity of H = 2.1)
(1)1.64
(2)1.82
(3)2.58
(4)2.91
can u explain the the solution..please..
1
Expert's answer
2012-10-17T08:50:06-0400
Xc-Xh= (eV)^-1/2 (Ed(HC)-[Ed(CC)+Ed(HH)]/2)^-1/2

1
kJ·mol−1 = 1.04×10-2 eV

Xc-Xh= (eV)^-1/2* (412-(348+436)/2)^-1/2 =4.47
kJ=(eV)^-1/2 * 4.47 *1.04×10-2 eV=0.22

Xc=2.1+0.22=2.32 it is closer to
(3) 2.58